### v1

# Semi-Supervised In-Context Learning: A Baseline Study

Abstract
Most existing work in data selection for In-Context Learning (ICL) has focused on constructing demonstrations from ground truth annotations, with limited attention given to selecting reliable self-generated annotations. In this work, we propose a three-step semi-supervised ICL framework: annotation generation, demonstration selection, and semi-supervised inference. Our baseline, Naive-SemiICL, which prompts select high-confidence self-generated demonstrations for ICL prompting, outperforms a 16-shot baseline by an average of 9.94% across 16 datasets. We further introduce IterPSD, an annotation approach that refines pseudo-demonstrations iteratively, achieving up to 6.8% additional gains in classification tasks. Lastly, we reveal a scaling law for semi-supervised ICL, where models achieve optimal performance with over 1,000 demonstrations. Code is available at: https://anonymous.4open.science/r/semi-supervised-icl-FA07
[…]
Semi-Supervised In-Context Learning: A Baseline Study
Zhengyao Gu1thanks: Equal Contribution., Henry Peng Zou [^1] footnotemark: 1, Yankai Chen2, Aiwei Liu3, Weizhi Zhang1, Philip S. Yu1 1University of Illinois Chicago  2Cornell University  3Tsinghua University {zgu24,pzou3,wzhan42,psyu}@uic.edu yankaichen@acm.orgliuaw20@mails.tsinghua.edu.cn
1 Introduction
In-Context Learning (ICL) has emerged as a powerful paradigm in natural language processing, enabling Language Models (LMs) to learn, adapt, and generalize from examples presented within their input context. This approach eliminates the need for extensive retraining and parameter modifications, facilitating more flexible and efficient learning Brown et al. (2020); Min et al. (2022); Agarwal et al. (2024); Fang et al. (2025). The effectiveness of ICL is contingent on the quality of the annotated examples, which encompasses factors such as diversity He et al. (2023), representativeness Yang et al. (2023), and ordering Lu et al. (2021). While significant efforts have been dedicated to selecting high-fidelity demonstrations using ground-truth annotations, the selection of reliable self-generated annotations for ICL remains underexplored. Existing methods are often constrained to specific tasks Zhang et al. (2023); Mamooler et al. (2024); Li et al. (2024a) and rely on task-specific heuristics Zhang et al. (2023). For instance, Auto-CoT Zhang et al. (2023) enhances demonstration quality by ensuring diversity and applying heuristic-based filtering rules, such as constraints on the size of reasoning steps. PICLe Mamooler et al. (2024) leverages LLM self-verification to discard demonstrations with incorrect named entity types, while SAIL Li et al. (2024a) filters out predictions that fail to translate back into the original input.
Figure 1: Semi-supervised ICL Framework. Ground truth data are used as demonstration for generating pseudo-demonstrations from unannotated data. The generated pseudo-demonstrations conjunctively with a small ground truth demonstration, are selectively used as demonstrations for the final prompting.
To advance domain-agnostic approaches for self-annotated demonstration selection, we propose a three-step framework (Figure 1) comprising: (1) annotation generation, (2) demonstration selection, and (3) semi-supervised inference. Drawing inspiration from semi-supervised learning, we term this framework Semi-Supervised ICL. We introduce a simple baseline, Naive-SemiICL, which selects high-confidence self-generated annotations as demonstration examples. Further extending this analogy, we propose IterPSD, an iterative approach for generating pseudo-demonstrations that progressively refines prompts by incorporating self-generated annotations at each step.
We systematically evaluate Naive-SemiICL against a robust 16-shot baseline using three confidence scoring methods across 16 datasets spanning 9 tasks. Experimental results indicate that Naive-SemiICL consistently outperforms the 16-shot baseline, achieving an average performance improvement of 9.94% on GPT-4o-mini (Figure 2). Furthermore, we demonstrate that IterPSD enhances semi-supervised ICL performance in five classification tasks, yielding gains of up to 6.8% (Table 1). Additionally, we identify a scaling law in semi-supervised ICL, where inference accuracy improves with increasing pseudo-demonstration examples, reaching optimal performance beyond 1,000 examples (Figure 3).
Figure 2: Comparison of GPT-4o-mini (top) and GPT-4o (bottom) performance across multiple datasets using three different methods: Few-Shot, Naive-SemiICL (Naive), and Naive-SemiICL without data selection (Unfiltered).
2 Method
In this section, we establish the framework of Semi-Supervised ICL, which consists of three phases: ① pseudo-demonstration generation, ② demonstration selection and ③ semi-supervised inference. Then, we propose a simple baseline for semi-supervised ICL, Naive-SemiICL, that generates pseudo-demonstrations in one iteration and filters out examples with a low confidence score. Building on top of Naive-SemiICL, we propose an iterative method for generating pseudo-demonstrations, IterPSD, that progressively enhances its prompt by incorporating self-generated annotations during the demonstration generation process.
2.1 Semi-Supervised ICL
Confidence-Aware In-Context Learning extends traditional ICL by outputting an additional confidence score for each input:
$(y,r,c)=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E},x)$ (1)
Like traditional ICL, the LLM is prompted with a task instruction $\rho_{\mathcal{T}}$ associated with task $\mathcal{T}$, a set of demonstrations $\mathcal{E}$, and an input $x$. Unlike traditional ICL, however, the model additionally returns a confidence score $c$ along with the predicted output $y$ and rationale $r$ [^1] In practice, generating rationales is optional. For example, one can query the LLM to directly generate the answer to a mathematical problem without intermediate reasoning steps., providing a measure of certainty for its predictions. We discuss the specific choice of confidence measure in Section 3.
Semi-Supervised ICL. Beyond ground-truth annotations, Semi-Supervised ICL leverages un-annotated data to enrich demonstrations. The setting assumes the availability of a ground-truth dataset $\mathcal{D}_{g}=\{(x_{i},y_{i})\}^{N_{l}}$, usually small in quantity, alongside a large pool of un-annotated data $\mathcal{X}_{u}={x_{i}}^{N_{u}}$. Semi-Supervised ICL augments the limited pool of ground-truth examples by generating pseudo-demonstrations $\mathcal{D}_{\mathrm{PSD}}$ from the un-annotated data. Formally,
$\mathcal{D}_{\mathrm{PSD}}=\{(x,\tilde{r},\tilde{y},\tilde{c})|x\in\mathcal{X}_{u}\},$ (2)
where $(\tilde{y},\tilde{r},\tilde{c})=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E}_{g},x)$ are generated by the LLM in an ICL fashion using a set of ground-truth demonstrated ns $\mathcal{E}_{g}\subseteq\mathcal{D}_{g}$. Low-confidence examples are then filtered out according to a confidence threshold $\lambda$:
$\mathcal{D}_{\mathrm{PSD}}^{\lambda}=\{(x,\tilde{r},\tilde{y},\tilde{c})|\tilde{c}\geq\lambda,(x,\tilde{r},\tilde{y},\tilde{c})\in\mathcal{D}_{\mathrm{PSL}}\}.$ (3)
During Semi-Supervised ICL inference, we sample pseudo-demonstrations $\mathcal{E}_{\mathrm{PSD}}$ from the filtered set $\mathcal{D}_{\mathrm{PSD}}^{\lambda}$, which we detail the specific methods in Appendix B.5. The LLM is then prompted with these pseudo-demonstrations alongside the ground-truth demonstrations from which they were generated:
$(\hat{y},\hat{r},\hat{c})=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E}_{l}\cup\mathcal{D}_{\mathrm{PSD}}^{\lambda},x).$ (4)
Most of the internal mechanisms of Semi-Supervised ICL are encapsulated by Equation 2, where pseudo-demonstrations are generated, while Equation 3 and Equation 4 represent simple operations. Next, we introduce two approaches for generating pseudo-demonstrations.
Figure 3: Pseudo-demonstration scaling (PSD) performance of Naive-SemiICL on GPT-4o (top) and GPT-4o-mini (bottom) across multiple datasets. The blue solid line represents Naive-SemiICL performance, and the dashed gray line represents the few-shot baseline.
2.2 A Simple Semi-Supervised ICL Baseline
We proposed a simple method that generates the pseudo-demonstrations in one iteration (Algorithm 1). We dub the method, together with the rest of the Semi-Supervised ICL framework, Naive-SemiICL. The method simply goes through the un-annotated data and generates a prediction, a rationale and a confidence score for each input. As the simplest form of Semi-Supervised ICL, it provides effective in-context learning signals by filtering out low-quality pseudo-demonstrations. We experiment with three commonly used confidence measures on 15 datasets spanning 9 tasks and show that this simple baseline consistently outperforms a strong 16-shot ICL baseline (Section 4.1).
Algorithm 1 Naive-SemiICL.
1:  Input: prompt $\rho_{\mathcal{T}}$, ground-truth demonstrations $\mathcal{E}_{l}\subseteq\mathcal{D}_{l}$, confidence score $\mathcal{C}$;
2:  Initialize $\mathcal{D}_{\mathrm{PSD}}=\emptyset$;
3:  for $x\in\mathcal{X}_{u}$ do
4:     $\tilde{y},\tilde{r},\tilde{c}=\mathrm{LM}(\rho_{\mathcal{T}},\mathcal{E}_{l},x)$;
5:     $\mathcal{D}_{\mathrm{PSD}}=\mathcal{D}_{\mathrm{PSD}}\cup\{(x,\tilde{r},\tilde{y},\tilde{c})\}$;
6:  end for
7:  Return $\mathcal{D}_{\mathrm{PSD}}$;
2.3 Iterative Pseudo-Demonstration Generation
Encouraged by the success of Naive-SemiICL, we explore whether pseudo-demonstrations can enhance the accuracy of subsequent pseudo-demonstration generation. We propose IterPSD (Algorithm 2), an iterative method for generating pseudo-demonstrations that
  1. 1.
    recursively adds newly generated pseudo-demonstrations to its own prompt before reaching maximum number of allowed demonstrations (Line 6), and
  2. 2.
    re-sample the most confident pseudo-demonstrations according to a confidence threshold $\lambda$ from all previously annotated instances once the size of the demonstration reaches a limit (Line 12).
In each iteration, IterPSD samples and annotates $K$ un-annotated examples before applying a filtering step. The generated pseudo-demonstrations are recursively accumulated and fed back into the LLM to generate additional pseudo-demonstrations (Line 10). To mitigate performance degradation caused by long context lengths, we impose an upper limit $\kappa$ on the number of self-fed pseudo-demonstrations. Once this limit is reached, the $\kappa$ most confident pseudo-demonstrations are resampled from the entire generated set, ensuring that only high-quality examples are retained (Line 7).
Order of Annotation.
To enhance the accuracy of pseudo-demonstration annotations, we introduce the $\epsilon$-Random Sampler (Algorithm 3), a sampling strategy that selects both similar and diverse examples from the un-annotated pool. At each iteration, a proportion $(1-\epsilon)$ of examples is chosen based on their cosine similarity to the nearest previously annotated instances (Line 6) [^2] We compute cosine similarity using vector embeddings generated by OpenAI’s text-embedding-3-large., ensuring that each selected example is closely related to an existing annotation. The underlying intuition is that the model achieves higher annotation accuracy on similar examples, thereby improving the quality of subsequent annotations. This approach aligns with curriculum learning Soviany et al. (2021), a widely used technique in semi-supervised learning that prioritizes simpler, well-understood examples before progressively annotating more challenging ones.
Algorithm 2 IterPSD
1:  Input: prompt $\rho_{\mathcal{T}}$, ground-truth demonstrations $\mathcal{E}_{l}\subseteq\mathcal{D}_{l}$, chunk size $K$, ratio of random examples $\epsilon$, maximum number of pseudo-demonstrations $\kappa$, confidence score $\mathcal{C}$;
2:  Initialize $\mathcal{D}_{\mathrm{PSD}}=\emptyset$; {Set of all the annotated pseudo-demonstrations.}
3:  Initialize $\mathcal{D}_{\mathrm{PSD}}^{\lambda}=\emptyset$
4:  Initialize $\overline{\mathcal{D}}_{\mathrm{PSD}}=\mathcal{X}_{u}$; {Set of un-annotated data yet to be annotated.}
5:  while $\overline{\mathcal{D}}_{\mathrm{PSD}}\neq\emptyset$ do
6:     if $|\mathcal{D}_{\mathrm{PSD}}^{\lambda}|>\kappa$ then
7:        $\mathcal{D}_{\mathrm{PSD}}^{\lambda}=$ top-$\kappa$ confident examples in $\mathcal{D}_{\mathrm{PSD}}$;
8:     end if {Cap the demonstration at a maximum size, prevent performance degradation from long-context.}
9:     $S=\mathrm{Sampler}(\mathcal{E}_{\mathrm{Iter}},\overline{\mathcal{D}}_{\mathrm{PSD}},K,\epsilon)$; {Retrieves a sample of size $K$ using $\epsilon$-Random Sampler}
10:     $\mathcal{D}_{S}=\text{Naive-SemiICL}(S,\rho_{\mathcal{T}},\mathcal{E}_{l}\cup\mathcal{D}_{\mathrm{PSD}}^{\lambda})$; {One iteration of Naive-SemiICL.}
11:     $\mathcal{D}^{\lambda}_{S}=\{(x,\tilde{r},\tilde{y},\tilde{c})|(x,\tilde{r},\tilde{y},\tilde{c})\in\mathcal{D}_{S},\tilde{c}\geq\lambda\}$; {Filter by confidence.}
12:     $\mathcal{D}_{\mathrm{PSD}}^{\lambda}=\mathcal{D}_{\mathrm{PSD}}^{\lambda}\cup\mathcal{D}^{\lambda}_{S}$;
13:     $\mathcal{D}_{\mathrm{PSD}}=\mathcal{D}_{\mathrm{PSD}}\cup\mathcal{D}_{S}$;
14:     $\overline{\mathcal{D}}_{\mathrm{PSD}}=\overline{\mathcal{D}}_{\mathrm{PSD}}-\mathcal{D}_{S}$;
15:  end while
16:  Return $\mathcal{D}_{\mathrm{PSD}}$;
Mitigating Confirmation Bias.
To maintain annotation quality, we ensure that at least half of the data ($\epsilon\geq 0.5$) is sampled randomly (Line 7) [^3] We found best performing $\epsilon$ to be $0.8$ in our experiments.. Without this, annotation accuracy degrades significantly. When $\epsilon=0$, selections are exclusively based on similarity to previously annotated examples. Pseudo-demonstrations become overly homogeneous, leading to bias in ICL predictions. This phenomenon closely parallels confirmation bias in semi-supervised learning Arazo et al. (2019); Zou and Caragea (2023), highlighting a strong connection between Semi-Supervised ICL and traditional semi-supervised learning frameworks.
Algorithm 3 $\epsilon$-Random Sampler
1:  Input: annotated demonstration $\mathcal{D}_{l}$, un-annotated demonstration $\overline{\mathcal{D}}_{l}$, chunk size $K$, random ratio $\epsilon$, prompt $\rho_{\mathcal{T}}$, embedder $\phi$.
2:  Initialize $S=\emptyset$;
3:  $K_{\mathrm{random}}=\epsilon K,K_{\mathrm{sim}}=(1-\epsilon)K$;
4:  Compute $d_{ij}=\mathrm{sim}_{\mathrm{cos}}(\phi(x_{i}),\phi(x_{j}))$ for all $x_{i}\in\mathcal{D}_{l},x_{j}\in\overline{\mathcal{D}}_{l}$;
5:  Compute $d_{j}=\min_{i}d_{ij}$ for all $x_{j}\in\overline{\mathcal{D}}_{l}$; {Compute distance to the nearest annotated example.}
6:  $S_{\mathrm{sim}}=\{x_{j}|d_{j}\in\text{Smallest}_{K_{\mathrm{sim}}}\{d_{j}\}\}$; {select the $K_{\mathrm{sim}}$ examples with the smallest distance to its nearest annotated demonstrations}
7:  Compute $S_{\mathrm{random}}$, a random sample of size $K_{\mathrm{random}}$ from $\overline{\mathcal{D}}_{l}-S_{\mathrm{sim}})$;
8:  $S=S_{\mathrm{sim}}\cup S_{\mathrm{random}}$;
9:  Return $S$;
3 Experimetnal Setup
Tasks and Datasets.
Our evaluation consists of 16 datasets spanning 9 tasks and 3 task types:
  • Classification. We include BANKING77 Casanueva et al. (2020) and CLINC Larson et al. (2019) for intent classification, CLINC(D) for domain classification, FewEvent Deng et al. (2020) for few-shot event classification, and FP Malo et al. (2013) for sentiment analysis.
  • Translation. We evaluate Naive-SemiICL’s ability to translate English into low-resource language using the FLORES200 dataset Costa-Jussà et al. (2022). In particular, we evaluate on six languages: Bemba, Fijian, Faroese, Tuvan, Venetian, and Sardinian.
  • Reasoning. We evaluate our method’s reasoning skills on GPQA Rein et al. (2024), a graduate-level scientific reasoning task, LiveBench Math White et al. (2025), a competition-level mathematics problems dataset, and three tasks from BigBenchHard Suzgun et al. (2022): Logical7, which tests LLM’s spatial reasoning ability, Geometric Shapes, which tests ability to interpret and identify geometric figures, and Date, which evaluates time manipulation skills.
We detail how we split these datasets for demonstration and evaluation in Appendix B.1.
Evaluation Metrics. For all classification and reasoning tasks, we report accuracy as the performance metric. We evaluate the equivalence of latex-style math outputs on LiveBench Math using the parser reported in Gao et al. (2024). For translation tasks, we report ChrF++ score Popović (2015) in its default configuration as implemented by TorchMetricsDetlefsen et al. (2022) following Agarwal et al. (2024). We report the average and standard error of three trials OpenAI (2024) on baseline results (Section 4.1). The rest of the results are collected over one trial.
Baselines. On classification and translation tasks, we compare Naive-SemiICL to a $16$-shot baseline. We report the best performing $k$-shot baseline where $k\leq 16$ on reasoning tasks, as the best-performing baseline does not necessarily take 16 demonstrations (see details in Appendix B.3). During pseudo-demonstration generation, we use the same ground truth demonstration examples that the baseline uses. We generate the label/translation directly on classification and translation tasks and leverage Chain-of-Thought prompting to generate rationale on reasoning tasks (see Appendix A for the CoT prompt). We experiment with different pseudo-demonstration sizes $k_{u}$ for Naive-SemiICL: $k_{u}\in\{32,64,100,150,200\}$ on classification tasks, and $k_{u}\in\{32,64,100,150,200\}$ on translation tasks.
Confidence Scores.
We primarily evaluate three confidence metrics: Verbalized Confidence, Entropy, and Self-Consistency, as defined below.
Verbalized Confidence. Verbalized Confidence Xiong et al. (2024) prompts the LLM to generate the confidence score as part of its response. See Appendix A for how we induce the Verbalized Confidence scores from LLMs in prompts.
Entropy. Entropy Shannon (1948) estimates the uncertainty of generated content using the token probability $P(w_{i}\mid w_{<i})$, where $w_{i}$ denotes the generated tokens and $w_{<i}$ represents the preceding tokens in the prompt:
$c_{\mathrm{Ent}}=-\frac{1}{L}\sum_{i=s}^{L}\log P(w_{i}\mid w_{<i}).$ (5)
We find Entropy unsuitable for estimating uncertainty on classification tasks, as it predominantly returns a confidence score of one, making the data selection step redundant.
Method BANKING CLINC CLINC(D) FewEvent FP
Naive-V 75.67 69.00 90.00 66.50 98.00
Naive-S 75.00 73.50 91.50 69.00 98.00
Iter-V 78.00 69.00 90.50 73.50 98.00
Iter-S 78.00 78.50 94.50 70.00 98.50
Improvement 3.10% 6.80% 3.28% 6.52% 0.50%
Table 1: Comparison of Naive-SemiICL (Naive) and IterPSD (Iter) methods on various datasets using GPT-4o-mini, evaluated using verbalized (-V) and self-consistency (-S) confidence scores. The best-performing results for each dataset are highlighted in bold, while the second-best results are underlined.
Self-Consistency. Self-Consistency Wang et al. (2023) samples multiple responses using diverse decoding paths and selects the most consistent answer based on majority voting. The relative frequency of the majority answer naturally defines a confidence score for the generated annotations. Specifically, let $K_{i}$ be the number of times the response $y_{i}$ appears in the set of $K$ sampled responses. The Self-Consistency confidence score is then computed as:
$c_{\mathrm{SC}}=\max_{i}\frac{K_{i}}{K}.$ (6)
We discuss how the confidence scores are applied to each task in Appendix B.4.
Hyperparameters.
Unless stated otherwise, we filter all generated pseudo-demonstrations by the confidence at the 90th percentile. We discuss the hyperparameters of IterPSD in Appendix B.7.
Models.
We experiment with GPT-4o-mini and GPT-4o checkpointed on 2024-07-18 and 2024-11-20 on all of our experiments.
4 Empirical Analyses
4.1 Naive-SemiICL Consistently Beat Few-Shot Baselines
Figure 2 presents a comparison of the Few-Shot baseline and Naive-SemiICL (Naive). For Naive-SemiICL, We report the performance of the best performing confidence score [^4] A detailed breakdown of the performance across different confidence scores is provided in Appendix C.. Naive-SemiICL outperforms the baseline on all tasks, except on CLINC(D) under GPT-4o, where it ties with the baseline. To assess the impact of demonstration selection, we also examine the performance of Naive-SemiICL without the filtering step (Unfiltered). The unfiltered setting yields lower performance compared to Naive-SemiICL on all task except on CLINC(D) and FP under GPT-4o, and performs worse than the baseline in 20 out of 32 settings, which indicates that pseudo-demonstrations alone are insufficient to guarantee improvement over the baseline.
Overall, GPT-4o-mini benefits more from Naive-SemiICL, achieving a 9.94% performance gain over the baseline, while GPT-4o sees only a 6.05% improvement. This disparity may be attributed to GPT-4o’s stronger few-shot capabilities, as it outperforms GPT-4o-mini by an average of 4.93% in few-shot settings across all tasks.
4.2 IterPSD Improves Upon Naive-SemiICL
Table 1 demonstrates the superiority of IterPSD over Naive-SemiICL across five classification tasks using GPT-4o-mini. We evaluate both methods with Verbalized Confidence and Self-Consistency, with the best-performing hyperparameters detailed in Appendix B.7. IterPSD consistently outperforms Naive-SemiICL, particularly when paired with Self-Consistency, though the magnitude of improvement varies by dataset. Notably, IterPSD (Self-Consistency) achieves significant gains on BANKING, CLINC, CLINC(D) and FewEvent (over 3.0% performance gain), except for FP.
Figure 4: We compare Naive-SemiICL accuracy across different ground truth demonstration sizes, with baseline performances indicated by dashed lines. On FewEvent, the maximum number of pseudo-demonstrations is capped at 1000 due to the limited availability of pseudo-demonstrations after filtering.
4.3 Scaling Law for Semi-Supervised ICL
Agarwal et al. (2024) identified a scaling law for In-Context Learning, where LLMs achieve peak performance with hundreds or thousands of ground-truth demonstrations. We observe a similar scaling law for Semi-Supervised ICL across classification and translation tasks. Figure 3 illustrates performance trends as a function of the number of sampled pseudo-demonstrations.
Across all configurations, Naive-SemiICL improves with larger demonstration sizes, though the point of peak performance varies. GPT-4o-mini scales effectively across most tasks, typically peaking between 500 and 1,000 examples, except for FP (100 examples) and Venetian (64 examples), where performance plateaus earlier. Similarly, GPT-4o reaches peak performance within the 500–1,000 range for classification tasks, with BANKING77 as an exception, achieving optimal performance with just 100 demonstrations. On translation tasks, GPT-4o exhibits a more stable scaling trend than GPT-4o-mini, with performance peaking later and declining more gradually.
We hypothesize that Naive-SemiICL’s decline in performance beyond a certain demonstration size stems from the accumulation of error in pseudo-demonstrations. To isolate the negative impact of long contexts on the LLMs, we examine the scaling behavior when all demonstrations are ground-truth data. Figure 5 presents results on six representative datasets, showing that both GPT-4o-mini and GPT-4o continue to improve beyond the peak of Naive-SemiICL. This suggests that the primary bottleneck lies in the quality of pseudo-demonstrations rather than the number of demonstrations.
4.4 Expanded Ground Truth Data
Is Semi-Supervised ICL still effective as the amount of ground-truth data increases, or does its effectiveness diminish with greater data availability? To answer this question, we evaluate Naive-SemiICL’s performance when $k_{g}\in\{32,64,100,500\}$ ground-truth demonstrations are available. We report the results using GPT-4o-mini with Verbalized Confidence on BANKING77, FewEvent and Bemba in Figure 4.
Across all three datasets, Naive-SemiICL outperforms most baselines, except for $k_{g}=100$ in FewEvent. The best performances are achieved with $k_{g}=500$, which performs on par with $k_{g}=64$ on FewEvent. We observe diminishing returns in performance gains as the number of annotated demonstrations increases. $k_{g}=64$ improves performance by 10.49% over the baseline, whereas $k_{g}=500$ yields only a 4.73% improvement on average across the three tasks. This indicates that increasing the number of ground-truth demonstrations improves performance, albeit with reduced gains. Additionally, we observe a scaling trend in BANKING77 and FewEvent similar to that reported in Section 4.3, but a generally downward trend on Bemba.
Figure 5: Many-shot scaling trends on classification and translation datasets. The blue solid line represents the performance of GPT-4o and GPT-4o-mini, while the gray dashed line indicates the number of demonstrations at which Naive-SemiICL achieves peak performance.
5 Related Work
Self-Generated Demonstrations.
Large Language Models (LLMs) exhibit remarkable zero-shot capabilities, allowing them to perform tasks without task-specific fine-tuning or prior examples Kojima et al. (2022). Their zero-shot predictions have proven to be effective sources of demonstration for in-context learning.
Auto-CoT Zhang et al. (2023) prompts the LLM with self-generated rationales on diversely sampled inputs. Rationales consisting of more than five reasoning steps are excluded from the demonstration to maintain the simplicity and accuracy of the demonstration. Such task-specific heuristic does not generalize to most recently published datasets such as LiveBench Math, as most of the generated rationale contains more than five steps. Our proposed Verbalized Confidence and Self-Consistency methods are task-agnostic and generally effective using the same cutoff threshold. Like Auto-CoT, Reinforced ICL Agarwal et al. (2024) generates rationales for reasoning problems and filters out those leading to incorrect answers. While this method requires the ground truth answer, our filtering method selects high-quality demonstration solely by the confidence score.
PICLe Mamooler et al. (2024) generates new demonstrations by annotating unlabeled examples and filtering out those with incorrect named entity types through self-verification prompting. Similarly, SAIL Li et al. (2024a) employs an annotation strategy for the bilingual lexical induction task, discarding predictions that fail to translate back to the original input. Both methods rely on task-specific filtering and require additional LLM queries for self-verification or back-translation. In contrast, our Verbalized Confidence approach is task-agnostic and requires only a single prompt for pseudo-labeling, significantly reducing inference overhead. Z-ICL Li et al. (2024b) leverages the zero-shot generative capability of large language models to synthesize demonstrations for subsequent in-context learning inference. In contrast, our approach assumes access to abundant unlabeled data and a small set of ground-truth labels, using the LLM only for annotation rather than for input generation.
Many-Shot ICL.
Agarwal et al. (2024) observed a significant performance increase in a variety of generative and discriminative tasks, as well as a scaling law between the number of examples in the demonstration and ICL performance. Our method hinges on this ability as our proposed method, Naive-SemiICL, fits at least 64 examples in the prompt. We report a similar scaling law for Semi-Supervised ICL in this work.
Traditional Semi-Supervised Learning.
Semi-supervised learning seeks to reduce reliance on labeled data by leveraging abundant unlabeled data to enhance model performance Lee et al. (2013); Sohn et al. (2020); Zou et al. (2025). Self-training McLachlan (1975); Xie et al. (2020) iteratively refines the model by using its own predictions on unlabeled data for training. Pseudo-labeling Lee et al. (2013); Sohn et al. (2020); Zou et al. (2023a, b) employs confidence-based filtering, retaining only high-confidence pseudo-labels to reduce error propagation and confirmation bias. JointMatch Zou and Caragea (2023) further alleviates error accumulation by using two independently initialized networks that teach each other through cross-labeling. Our work is the first to integrate confidence filtering and leverage both labeled and pseudo-labeled data in an in-context learning framework.
6 Conclusion
We introduced a semi-supervised In-Context Learning (ICL) framework that enhances self-generated pseudo-demonstrations through confidence-based data selection and iterative annotation. Our simple semi-supervised method, Naive-SemiICL, outperformed a strong 16-shot ICL baseline, achieving an average performance gain of 9.94% across 16 datasets. We further proposed IterPSD, an iterative refinement approach for pseudo-demonstrations, yielding up to 6.8% additional gains in classification tasks. Moreover, we identified a scaling law in semi-supervised ICL, showing that models achieve optimal performance with over 1,000 pseudo-demonstrations. Our analysis of Naive-SemiICL with increasing ground-truth data revealed diminishing returns—while additional ground-truth annotations improve performance, the relative contribution of pseudo-demonstrations decreases. This suggests that semi-supervised ICL is effective in low-resource settings, yet remains beneficial even with larger ground-truth datasets.
7 Limitation
While this work investigates the potential of semi-supervised ICL, several limitations remain. First, the reliance on SOTA LLMs for ICL introduces substantial computational overhead, posing challenges for researchers and practitioners with limited resources. Second, although we have shown that the incorporation of pseudo-demonstration generation strategies enhances and improves ICL performance on two models, the effectiveness of our proposed method might be sensitive to the choice of model. Future work can explore more advanced confidence calibration techniques for pseudo-demonstration selection, such as adaptive thresholding. Additionally, noise-aware in-context learning remains an under-explored domain that could potentially improve the robustness of Semi-Supervised ICL.
Appendix A Prompts
The prompts are illustrated in Table 4. {CAPITAL LETTERS} enclosed in curly brackets are variables that are substituted during inference.
Appendix B Experimental Details
B.1 Train-Test Split
For classification tasks with more than 5,000 examples, we randomly sample 5,000 examples for demonstration and 200 for evaluation. For tasks with less than 5,000 examples, we randomly sample 200 for evaluation and use the rest for demonstration. Each FLORES dataset is comprised of a development set with 997 examples and a development test set with 1012 examples. We use all of 997 for demonstration and randomly sample 200 from the development test examples for evaluation. We use the diamond split (198 examples) of GPQA following Agarwal et al. (2024), out of which 99 are used for evaluation and the other 99 are used for demonstration. Since LiveBench Math contains math problems from three sources, we evenly sample 150 questions from different sources for evaluation and use the rest for demonstration. Each BigBenchHard dataset contains 250 examples. We randomly sample 100 for evaluation and use the rest for prompting.
B.2 Dataset Details
Classification Datasets.
  • BANKING77. The BANKING77Casanueva et al. (2020) dataset is a fine-grained intent classification benchmark in the banking domain, consisting of 13,083 customer queries labeled into 77 intent categories.
  • CLINC. The CLINC150 Larson et al. (2019) dataset is a benchmark for intent classification, containing 22,500 user queries across 150 intent categories grouped into 10 domains, along with an out-of-scope category. We refer to the intent classification task of CLINC150 as CLINC.
  • CLINC(D). We refer to the domain classification annotation of CLINC150 as CLINC(D).
  • FewEvent. The FewEventDeng et al. (2020) dataset contains 4,436 event mentions across 100 event types, with each event type having only a few annotated examples (typically 5 to 10 per type).
  • FP. Financial PhrasebankMalo et al. (2013) The Financial PhraseBank dataset consists of 4840 sentences from English language financial news categorised by sentiment.
Low-Resource Language Translation.
FLORES-200 Costa-Jussà et al. (2022) contains 200 languages translated from a common corpus. It is an extension of the original FLORES-101 Goyal et al. (2022) dataset, which covered 101 languages. The dataset covers low-resource and high-resource languages, including many languages with little prior data on. It includes many African, South Asian, and Indigenous languages, making it one of the most diverse multilingual benchmarks.
Reasoning Datasets.
  • GPQA. GPQARein et al. (2024) is a multiple-choice question answering benchmark, with graduate-level questions that involves reasoning in biology, physics, and chemistry.
  • LiveBench Math. LiveBenchMath contains 368 contamination-free mathematical problems, sampled from high school math competitions, proof-based fill-in-the-blank questions from Olympiad-level problems, and an enhanced version of the AMPS dataset.
  • BigBenchHard. We include three tasks from BigBenchHardSuzgun et al. (2022). Logical7 evaluates a model’s ability to deduce the order of a sequence of objects based on provided clues about their spatial relationships and placements. The Geometric Shapes task within the BigBenchHard evaluates a model’s ability to interpret and identify geometric figures based on SVG path data. The Date task within the BigBenchHard benchmark evaluates a model’s ability to comprehend and manipulate date-related information.
GPQA LiveBench Math Logical7 Geometric Shapes Date
GPT-4o-mini 4 0 8 8 0
GPT-4o 4 0 8 8 16
Table 3: Best number of shots for the baseline on reasoning tasks.
B.3 Baseline Details
We observe that few-shot baselines (Section 3) does not necessarily scale with more demonstrations. Thus, we report the best performing $k$-shot baseline where $k\leq 16$, which we report in Table 3.
B.4 Applying Confidence Scores
On all tasks, we sample from the LLMs 10 times to compute the Self-Consistency score. Self-Consistency is unsuitable for translation tasks due to the computational challenges of assessing equivalence between translations. Instead, we introduce Back-Translation, which evaluates translation quality by translating the output back to the original language. The confidence score is then derived using the cosine similarity (on embeddings) between the back-translation and the original input. A detailed description of Back-Translation is provided in Appendix B.6.
B.5 Pseudo-Demonstration Sampling
On classification tasks, BigBenchHard tasks and GPQA, we sample diversely by evenly sample predicted labels. For translation tasks and LiveBench Math, we utilize OpenAI’s text-embedding-3-large to generate embeddings for each example input. Then we cluster the embeddings into clusters and evenly sample the most representative (the one closest to the cluster centroid) instances from each cluster.
B.6 Back-Translation
Suppose an LLM has translated a source language input $s$ into a target language output $t$. We then use the same LLM to translate $t$ back to the original language
$\hat{s}=\mathrm{LM}(t,\rho_{b}),$
where $\rho_{b}$ is a prompt that induces the back-translation. Then, the Back-Translation Confidence is the cosine similarity between the original input $s$ and the back-translation $\hat{s}$
$c=\mathrm{sim}_{\mathrm{cos}}(\phi(\hat{s}),\phi(s)),$
where $\phi$ is an embedding function.
B.7 Hyperparameters
Confidence Threshold.
We find that setting the confidence threshold $\lambda$ at the 90th percentile of all generate pseudo-demonstrations $\mathcal{D}_{\mathrm{PSD}}$ allowed Naive-SemiICL to achieve competitive performance. We assume this threshold in this work unless stated otherwise.
IterPSD
takes 3 hyperparameters: chunk size $K$, which controls how many data to annotate in each iteration, $\epsilon$, which controls the proportion of random sample in each chunk, $\kappa$, the maximum amount of examples allowed in the demonstration while generating pseudo-demonstrations. We experiment with $K\in\{100,500,1000\}$, $\epsilon\in\{0.5,0.8,1.0\}$, $\kappa\in\{300,500,1000\}$. We find that $\epsilon=0.8$, $K=500$, and $\kappa=1000$ yielded the best results on all tasks, except on FP, where $K=100$ and $\kappa=300$ yielded the best result.
Appendix C Effects of Different Confidence Methods
In this section, we examine the performance Naive-SemiICL paired with different confidence methods, which we compile as Table 2. We observe that classification and translation tasks each have a dominant confidence measure. For classification tasks, Self-Consistency emerges as the most effective confidence method. It surpasses the Verbalized Confidence method on 4 out of 5 datasets across both models. Verbalized Confidence is the leading measure for translation tasks, consistently achieving the highest performance across all languages. For reasoning tasks, no single method clearly dominates. Under GPT-4o-mini, Verbalized Confidence yields the best average performance, while under GPT-4o, Entropy slightly outperforms Self-Consistency, securing the top position by a narrow margin.
Overall, Self-Consistency improves classification and reasoning tasks, but its effect varies across translation tasks and is not applicable to all tasks. Entropy is sometimes useful in reasoning tasks, but fall short on translation tasks. Verbalized inference remains a strong and economical baseline across all tasks but is generally outperformed by Self-Consistency on classification tasks.
Table 4: The prompt template we use for classification, translation, and reasoning tasks, respectively.
Types Prompts
Classification You are a helpful assistant who is capable of performing a classification task (mapping an Input to a Label) with the following possible labels: {A LIST OF POSSIBLE LABELS} ___ Here are zero or more Input and Label pairs sampled from the classification task. {DEMONSTRATIONS} ___ Now, Label the following Input among the following Input: {INPUT}
[…]
Translation You are an expert translator. I am going to give you zero or more example pairs of text snippets where the first is in the source language and the second is a translation of the first snippet into the target language. The sentences will be written in the following format: <source language>: <first sentence> <target language>: <translated first sentence> ___ {DEMONSTRATIONS} ___ Now, Translate the following \$source text into \$target. Also give the Confidence of your given Answer in the following format: **Confidence**: <a confidence score between 0 and 1>: English: {INPUT SENTENCE} {TARGET LANGUAGE}:
[…]
Reasoning First, I am going to give you a series of Questions that are like the one you will be solving. ___ {DEMONSTRATIONS} ___ Now, Answer the following Question. Think step by step. Question: {QUESTION} Also give the Confidence of your given Answer in the following format: **Confidence**: <a confidence score between 0 and 1>